40×4g×40^3=bw^3E............(1)
60×4g×60^3=bw^3E............(2)
Solve for E and g simultaneously
1
Expert's answer
2021-08-11T09:54:07-0400

There is no solution:


"\frac{40\cdot4g\cdot40^3}{60\cdot4g\cdot60^3}=\frac{32}{81}\\
\frac{bw^3E}{bw^3E}=1"